# On the Birkhoff Spectrum for Hyperbolic Dynamics

### v1

@@ 3 Proof of Main Results for Diffeomorphisms
The proof of Theorem 3.7 will be a consequence of following Lemma, which is a consequence of proof of Theorem 1.1
@@
Now let us prove the density: If the interval $[m(f),M(f)]$ consists of a single point, i.e., $m(f)=M(f)$, the proof is trivial. We may therefore assume $m(f)<M(f)$.
Let $a,b$ be such that $m(f)<a<b<M(f)$. We will prove that there exists a periodic point $p\in\text{Per}(\varphi)$ for which the average $\frac{S(f,\varphi,p)}{\tau_{p}}$ lies in $(a,b)$.
Suppose, for contradiction, that no such periodic point exists. We then define the following two subsets of $\text{Per}(\varphi)$:
@@
Now, let $\beta=\frac{a+b}{2}$. By Theorem 1.1, the set $\mathcal{B}(f-\beta,\varphi,\Lambda)$ is dense in $\mathbb{R}$. Consider the set
@@
This set $\Delta_{\beta}$ must be infinite due to the density property.
We now examine the intersection of $\Delta_{\beta}$ with the sets defined above.
@@
    This implies $\tau_{p}<\frac{b}{b-\beta}$, so again the period is bounded. Hence, the set $\Pi_{b}\cap\Delta_{\beta}$ is also finite.
@@
By Theorem 1.1, we may assume without loss of generality that $\mathcal{B}(f,\varphi,\Lambda)\subset\mathbb{R}^{+}$. Under this assumption, we consider two cases:
@@
$\frac{S(f,\varphi,p)}{\tau_{p}}=\frac{S(f,\varphi,q)}{\tau_{q}}=\alpha\geq 0.$
Since $f$ is not cohomologous to 0, it follows that $\alpha>0$. Consequently,
@@
Choose a constant $\beta$ such that
$\frac{S(f,\varphi,p)}{\tau_{p}}<\beta<\frac{S(f,\varphi,q)}{\tau_{q}}$
and define the Hölder continuous function $g(x)=f(x)-\beta$. We then have:
$S(g,\varphi,p)=S(f,\varphi,p)-\beta\tau_{p}<0<S(f,\varphi,q)-\beta\tau_{q}=S(g,\varphi,q).$
By Theorem 1.1, the set $\mathcal{B}(g,\varphi,\Lambda)$ is dense in $\mathbb{R}$. Therefore, for any $M>0$, there exists a periodic point $p$ such that
@@ 3.4.1 Number-theoretic lemmas
Now, for the sake of contradiction, suppose $S\subseteq[A-\delta,A+\delta]$, for some $\delta<\dfrac{\gamma c}{4}$. This would imply $\mathrm{dist}(\frac{B_{n}}{c},\mathbb{Z})<\dfrac{\delta}{c}$ for all $n$. Which implies that $\dfrac{\gamma}{4}<\dfrac{\delta}{c}$. This implies a contradiction for the choice of $\delta_{0}=\dfrac{\gamma c}{4}$. It follows that for any $\delta<\delta_{0}$, the set $S$ cannot be a subset of $A-\delta,A+\delta]$, which is the desired result.
Lemma 3.14.
If $a_{1},a_{2},\dots,a_{l}\geq 0$, not all zero, then the set $\Big\{\sum_{i=1}^{l}\frac{n_{i}}{k}a_{i}:n_{i},k\in\mathbb{N}\Big\}$ is dense in $\mathbb{R}^{+}$.
Proof.
Put $\mathcal{A}:=\displaystyle\Big\{\sum_{i=1}^{l}\frac{n_{i}}{k}a_{i}:n_{i},k\in\mathbb{N}\Big\}$. Observe that $\inf\mathcal{A}=0$ Therefore, given $x\in\mathbb{R}^{+}$ and $\epsilon>0$ there is $a\in\mathcal{A}$ with $0<a<\epsilon$. Moreover, $x=na+r$ for some $n\in\mathbb{N}$ and $0\leq r<a$. So, $|x-na|=r<a<\epsilon$, the result follows since $na\in\mathcal{A}$. ∎
Lemma 3.15.
If $a,b\in\mathbb{R}$, then if $\frac{b}{a}\in\mathbb{Q}$, then the set $a\mathbb{N}-b\mathbb{N}$ is not dense.
To conclude this subsection, we prove Theorem 1.5. The proof leverages the hyperbolic structure of the basic set—using tools such as the shadowing lemma---to establish the appropriate setting for applying the preceding lemmas.
@@
We consider two cases based on the rationality of the ratios $\dfrac{a_{i}}{a_{j}}$ for $i\neq j$.
Case 1: All ratios $\frac{a_{i}}{a_{j}}$ are rational.
Proof of Case 1.
Suppose, for contradiction, that there exist periodic points $p,q\in\text{Per}(\varphi)$ satisfying
$\frac{S(f,\varphi,p)}{\tau_{p}}<\frac{S(f,\varphi,q)}{\tau_{q}}.$
By Lemma 3.14, there exist positive integers $n_{1},n_{2},\dots,n_{k}$ and $k\in\mathbb{N}$ such that
$\frac{S(f,\varphi,p)}{\tau_{p}}<\sum_{i=1}^{l}\frac{n_{i}}{k}a_{i}<\frac{S(f,\varphi,q)}{\tau_{q}}.$
Define $\beta:=\displaystyle\sum_{i=1}^{l}\frac{n_{i}}{k}a_{i}$. Theorem 1.1 then implies that $\mathcal{B}(f-\beta,\varphi,\Lambda)$ is dense in $\mathbb{R}$.
Observe that
$S(f-\beta,\varphi,p)\subset\displaystyle\bigcup_{i=1}^{l}a_{i}\mathbb{N}-\sum_{i=1}^{l}\frac{n_{i}}{k}a_{i}\mathbb{N}\subset\bigcup_{i=1}^{l}a_{i}\left(\left(1-\frac{n_{i}}{k}\right)\mathbb{N}-\sum_{j\neq i}\frac{n_{j}}{k}\frac{a_{j}}{a_{i}}\mathbb{N}\right).$
When $l=1$, we have $\beta=\frac{n_{1}}{k}a_{1}$ and consequently $S(f-\beta,\varphi,p)\subset a_{1}(\mathbb{N}-\frac{n_{1}}{k}\mathbb{N})$. However, Lemma 3.15 shows this set cannot be dense, contradicting Theorem 1.1.
We conclude that for all $p,q\in\text{Per}(\varphi)$,
$\frac{S(f,\varphi,p)}{\tau_{p}}=\frac{S(f,\varphi,q)}{\tau_{q}}=\alpha\quad\text{for some }\alpha\geq 0.$
Thus $\alpha=a_{i_{0}}$ for some $i_{0}\in\{1,\dots,l\}$, and Livšic’s Theorem completes the proof. ∎
Case 2: Some ratio $\frac{a_{i}}{a_{j}}$ is irrational.
Proof of Case 2.
Fix indices $i_{1},i_{2}$ such that $\frac{a_{i_{1}}}{a_{i_{2}}}\notin\mathbb{Q}$.
@@ Conclusion.
First for basic sets, assume $f$ is a dispersed function, meaning there exist $\theta,\omega\in\Lambda$ such that
@@
If $\phi^{t}$ Anosov, it is Axiom A and the result follow for the first case. For the case of transitive Anosov flows, If $f$ is dispersed, then there exist $\theta,\omega\in\operatorname{Per}(\phi^{t})$ such that

### v2

@@ 3.2 Example of dispersed and arithmetic Lipchitz Function on basic set
The proof of Theorem 3.6 will be a consequence of following two Lemmas.
Lemma 3.7.
Let $\Lambda$ be a basic set for a diffeomorphism $\varphi$, and let $f\colon\Lambda\to\mathbb{R}$ be continuous and not cohomologous to constant. For any interval $I\subset\mathbb{R}$, there exists $\beta\in I$ such that $f-\beta$ is non-arithmetic.
Proof.
If $p\in\operatorname{Per}(\varphi)$, set $A_{p}:=\dfrac{S(f,\varphi,p)}{\tau_{p}}$. Since $f$ is not cohomologous to a constant there are $p_{1}\neq p_{2}$ with $A_{p_{1}}\neq A_{p_{2}}$. For $\beta\neq A_{p_{2}}$ define
$R(\beta):=\frac{A_{p_{1}}-\beta}{A_{p_{2}}-\beta}.$
For each rational $\frac{m}{n}\in\mathbb{Q}\setminus\{0\}$, the equation $R(\beta)=\frac{m}{n}$ has at most one solution $\beta$. Hence
$B:=\{\,\beta\in I\mid R(\beta)\in\mathbb{Q}\,\}$
is countable. Pick $\beta\notin B\cup\{A_{p_{2}}\}$. Then $R(\beta)\notin\mathbb{Q}$.
Now, put
$u:=\tau_{p_{1}}(A_{p_{1}}-\beta)=S(f-\beta,\varphi,p_{1}),\qquad v:=\tau_{p_{2}}(A_{p_{2}}-\beta)=S(f-\beta,\varphi,p_{2}).$
Their ratio $\frac{u}{v}=\frac{\tau_{p_{1}}}{\tau_{p_{2}}}\,R(\beta)$, is irrational.
Assume, for a contradiction, that $f-\beta$ is arithmetic, then for some $c\neq 0$ the set of all such Birkhoff sums is contained in $c\mathbb{Z}$. Then $u=ck_{1}$ and $v=ck_{2}$ for integers $k_{1},k_{2}$, so $u/v=k_{1}/k_{2}\in\mathbb{Q}$, contradicting the irrationality of $u/v$.
Thus, no such $c$ exists, which proves the lemma. ∎
For the next lemma, we use the following Sigmund’s Theorem related to the density of periodic measures.
Theorem 3.8 ([17]).
If $\Lambda$ is a basic set for a diffeomorphism $\varphi$, then the set
Using Theorem 3.8 and Lemma 3.7, we have
@@
Now let us prove the density: If the interval $[m(f),M(f)]$ consists of a single point, i.e., $m(f)=M(f)$, the proof is trivial. We may therefore assume $m(f)<M(f)$ (in particular, $f$ is not cohomnologous to constant).
Let $a,b$ be such that $m(f)<a<b<M(f)$. We will prove that there exists a periodic point $p\in\text{Per}(\varphi)$ for which the average $\frac{S(f,\varphi,p)}{\tau_{p}}$ lies in $I:=(a,b)$.
Suppose, by contradiction, that no such periodic point exists. We then define the following two subsets of $\text{Per}(\varphi)$:
@@
For every $\beta\in I$ the function $f-\beta$ is dispersed. Moroever, from Lemma 3.7 there is $\beta\in I$ such that $f-\beta$ is non-arithmetic, then by Theorem 3.1, the set $\mathcal{B}(f-\beta,\varphi,\Lambda)$ is dense in $\mathbb{R}$. Consider the set
@@
By density, the set $\Delta_{\beta}$ is infinite due to the density property.
We now examine the intersection of $\Delta_{\beta}$ with the sets $\Gamma_{a}$ and $\Gamma_{b}$.
@@
    Since $b-\beta>0$, this implies $\tau_{p}<\frac{b}{b-\beta}$, so again we have a contradiction (if $b<0$) or the period $\tau_{p}$ is bounded (if $b>0$). Hence, the set $\Pi_{b}\cap\Delta_{\beta}$ is also finite.
@@
$\frac{S(f,\varphi,p)}{\tau_{p}}=\frac{S(f,\varphi,q)}{\tau_{q}}=\alpha.$
Since $f$ is not cohomologous to 0, it follows that $\alpha\neq 0$. Consequently,
@@
Consider the interval $I:=\Big(\frac{S(f,\varphi,p)}{\tau_{p}},\frac{S(f,\varphi,q)}{\tau_{q}}\Big)$. From Lemma 3.7, we can choose a constant $\beta\in I$ such that the Hölder continuous function $g(x)=f(x)-\beta$ is dispersed and non-arithmetic. Therefore, by Theorem 3.1, the set $\mathcal{B}(g,\varphi,\Lambda)$ is dense in $\mathbb{R}$. Therefore, for any $M>0$, there exists a periodic point $p$ such that
@@ 3.5.1 Number-theoretic lemmas
Now, for the sake of contradiction, suppose $S\subseteq[A-\delta,A+\delta]$, for some $\delta<\dfrac{\gamma c}{4}$. This would imply $\mathrm{dist}(\frac{B_{n}}{c},\mathbb{Z})<\dfrac{\delta}{c}$ for all $n$. Which implies that $\dfrac{\gamma}{4}<\dfrac{\delta}{c}$. This implies a contradiction for the choice of $\delta_{0}=\dfrac{\gamma c}{4}$. It follows that for any $\delta<\delta_{0}$, the set $S$ cannot be a subset of $A-\delta,A+\delta]$, which is the desired result. ∎
To conclude this subsection, we prove Theorem 1.3. The proof leverages the hyperbolic structure of the basic set—using tools such as the shadowing lemma---to establish the appropriate setting for applying the preceding lemmas. It is worth noting that the proof follows a similar overall approach to [6], but we provide full details here because we need to produce certain specific combinations of Birkhoff sums for our argument. Moreover, some of these combinatorial structures will be needed again in the proof of Theorem 1.2.
It is important to note, the proof also take advance in similar approch like [Shabao], but write all the details since we need some specfic combination/phenomenon. Moreover, some of this combinations will be necessary in the proof of Theorem 1.2
@@
First, if $f$ is dispersed, then by Theorem 3.1 and the fact that by hypotheses $f$ is non-arithmetic, the set $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $\mathbb{R}$. But $\mathcal{B}(f,\varphi,\Lambda)\subset\bigcup_{i}a_{i}\mathbb{Z}$ is a countable union of discrete additive subgroups, hence nowhere dense in $\mathbb{R}$; this contradicts density in $\mathbb{R}$. Therefore, $f$ cannot be dispersed. Consequently, we may assume $f$ is not dispersed, i.e.
$\mathcal{B}(f,\varphi,\Lambda)\subset\bigcup_{i=1}^{l}a_{i}\mathbb{N}.$
The proof of this theorem relies on the fact that for some $i\neq j$ the ratio $a_{i}/a_{j}$ is irrational.
Claim: There are $i\neq j$ such that $\dfrac{a_{i}}{a_{j}}\notin Q$.
Proof of Claim.
Assume, for the sake of contradiction, that all ratios $\dfrac{a_{i}}{a_{j}}$ are rational. Write each ratio $\dfrac{a_{i}}{a_{i+1}}$ in lowest terms as $\dfrac{p_{i}}{q_{i}}$, for $i=1,\dots l-1$, then
$\mathcal{B}(f,\varphi,\Lambda)\subset\bigcup_{i=1}^{l}a_{i}\mathbb{N}\subset\frac{a_{l}}{\prod_{i=1}^{l-1}q_{i}}\mathbb{N}$
Thus, we have a contradiction since $f$ is non-arithmetic. Hence the assumption that all ratios $\dfrac{a_{i}}{a_{j}}$ are rational is false, which completes the proof of the claim. ∎
From the Claim and the assumption that $f$ is non-arithmetic, we can assume that there are two different indices $i_{1},i_{2}$ such that $\dfrac{a_{i_{1}}}{a_{i_{2}}}\notin\mathbb{Q}$.
@@ Conclusion.
First for basic sets, assume $f$ is a dispersed and non-arithmetic function, meaning there exist $\theta,\omega\in\Lambda$ such that
@@
If $\phi^{t}$ Anosov, it is Axiom A and the result follow for the first case. For the case of transitive Anosov flows, If $f$ is dispersed and non-arithmetic, then there exist $\theta,\omega\in\operatorname{Per}(\phi^{t})$ such that
@@ 4.3 Some applications to geodesic flow
Acknowledgments: The author thanks Mingyang Xia for his invaluable insights, which were essential to the completion of this work.